Three sets of twins are to pose for photographs. How many photographs are possible if each set of the twins are to stand close to each other?

1
Expert's answer
2022-03-28T17:54:52-0400

Choose the order of pairs from left to right – there are "3!=6" ways. In each pair choose the order of twins – there are "2\cdot 2\cdot 2=8" ways. The total number of ways is "6\cdot 8=48"